OpenStudy (anonymous):

Find the slope of the line whose equation is 8y = 4 - 3x. -3/8 -8/3 2

OpenStudy (anonymous):

First, remember \[y=mx+b\] where m is the slope so right now you have \[8y=4-3x\] and you want \[y=f(x)\] or in otherwords you want y without the 8 or the constant

OpenStudy (anonymous):

y=(4-3x)/8 slope is y' that is the derivative of y y'=dy/dx y'=((4-3x)/8)' y'=-3/8

OpenStudy (anonymous):

derivatives are not needed for this

OpenStudy (anonymous):

if you divide by 8 and put it in the \[y=mx+b\] form, you'll see the slope

OpenStudy (anonymous):

yes, derivatives are not needed for such simple functions, but the generic approach is always more useful, as it can be used for all rational functions. For a simple function of type y=ax+b, the slope is simply a